The conversion of A reaches 50% in a five minute run. What is the order of the reaction if it takes 10 minutes to reach 75% conversion'! What is the order of the reaction if it takes 20 minutes to reach 75% conversion? What is the order of the reaction if it takes 30 minutes to reach 75% conversion?
Answer #1

solution t-bmin t = 10 min a? (a) XA = Ois XA, 075 (^&A)=-da akan at + => -1944 khain I CAM LAO inh - CAO 1 - И => -TA 16 J:to.bos (1-1) = 70-301 (tan) 2-20 1- и 1 = n. 1 nas (b) X=0.75 CAO In - (-15k na? t=20min from equation (1) Canon [11-0.75)*_,

From above this is clear that the Reaction is of first order.

Thankyou.... please upvote it ???